OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + 2y = 7 2x + y = 8

OpenStudy (anonymous):

x=7-2y therefore 14-4y+y=8 3y=6 y=2 thus x=3 (3,2)